Instrument Log  

C-Band Radar         IF7      Near Nardin, OK

 

Date

Technician

Comments

7/21/16

DKB

Site power was lost due to main power line cut by local farmer at 1700 GMT. Performed shutdown of the shelter and UPS’s at 1900 GMT.

7/20/16

DKB

Site visit performed weekly PM and found no problems at 1633 GMT.

7/13/16

DKB

Performed weekly radar inspections at 1546 GMT.

7/6/16

DKB

Performed weekly radar inspections at 1618 GMT.

6/29/16

DKB

Performed weekly radar inspections at 1823 GMT.

6/28/16

DKB

Found no new files on the web page since 6/28 0648 GMT. Found the PGEN eth0 graph stopped updating at 0645 GMT. Cycled the DAQ power to restart at 1314 GMT. Files for ppi returned but not for rhi. The transmitter power is also low. T/S in progress.

6/27/16

DKB

Found no new files on the web page since 6/27 0011 GMT. Found the PGEN eth0 graph stopped updating at 0001 GMT. Reset the Eng web page and cycled the DAQ power to restart at 1424 GMT.

6/24/16

DKB

After the system was re-started, noticed no new files on the web page. Performed web page reset and RTS at 2022 GMT.

6/24/16

BMA

Found the radar had a UPS fault while doing daily checks. Restarted PGEN and MGEN. Reset all I/O. Transmitter came back online. Started scan sequence.  RTS 6/24/16  1600 GMT.

6/15/16

DKB

Weekly site PM visit. RTS at 1715 GMT.

6/8/16

BMA

Weekly site PM visit. RTS at 1550 GMT.

5/25/16

BMA

Performed monthly instrument PM and inspections at 1815 GMT.

 

5/19/16

DKB

Performed weekly instrument PM and inspections at 1832 GMT.

 

5/12/16

BMA

The recently installed split unit HVAC PDU had a 20-30 min reaction time when turning on/off the outlet. Contacted vendor and they suggested a firmware update. Installed update and time delay problem was solved. The web interface still does not work as the Java software on the radar needs to be updated. Mentors are currently working that issue. RTS 5/12/16 2230 GMT.

5/11/16

DKB

Performed weekly instrument PM and inspections at 1943 GMT.

5/10/16

BMA

Transmitter would not operate. Mentor suggested we tune the magnetron IAW the operator handbook. Transmitter tuned correctly. RTS: 10 MAY 16 1920 GMT.

5/4/16

BMA

Performed weekly instrument PM and inspections at 1815 GMT.

5/2/16

BMA

Radar had a UPS fault and was not scanning. Went to site, reset all faults, and returned the radar to operation. RTS 5/2/16  1830 GMT.

4/28/16

BMA

Performed monthly instrument PM and inspections at 1850 GMT.

4/18/16

DKB

Found the system off line due to power outage overnight. Thunderstorms passed through the area. Reset all Mgen IO, reset the UPS alarms, restarted the scan, powered on the transmitter and RTS 1510 GMT.

4/6/16

BMA

Mentors requested PDUs be installed to control HVAC systems. PDUs installed. Shutdown: 4/6/16 1400 GMT; RTS: 4/6/16 1830 GMT

3/30/16

BMA

Performed the monthly instrument PM and inspections at 1428 GMT.

3/23/16

DKB

Weekly PM no problems noted. RTS 2145 GMT.

3/15/16

BMA

When doing checks, noticed that Transmitter Power in BITE software read 0. Reset RCM. Power reading intermittently comes and goes but is consistently correct when read through terminal commands. Mentor suggested it could be a problem with the BITE software and that we continue to monitor it. Arrived onsite: 1800 GMT, RTS: 1900 GMT.

3/7/16

BMA

Performed the monthly instrument PM and inspections at 1706 GMT.

2/24/16

BMA

Performed the monthly instrument PM and inspections at 1736 GMT.

2/22/16

BA

When SGP techs shut down radar for checks, the incorrect scans were started when the system was brought back up.  Programmed correct scans through radar GUI.

2/19/16

BA

Camera in radome was inop.  Tested and reseated power and network cables. Camera came back online.

2/3/16

BA

Cover for pedestal power plate missing.  Fabricated and replaced cover.

1/26/16

BA

Radar Az motor was leaking oil.  Kim Weaver from ARC came out and assisted with replacing the *Az motor.

*Old – 2300813   *New - 2856001 WD80034

12/1/15

BA/CM

Radar return power was low.  Adjusted magnetron frequency from 5.623 GHz to 5.625 GHz at 2014 GMT.  Also found that one of the LNAs was loose.  Tightened LNA SMA connections.

11/6/15

BA

CSAPR was taken down at 1500 GMT to test spare Az motor.  Motor was tested and system was RTS at 1630 GMT.

10/20/15

BA

UPS had a power fault due to possible power loss.  Reset all faults and restarted system at 1930 GMT.

9/24/15

BA

PGEN computer was not loading right bios on power up.  Mentors reviewed bios and and reloaded to correct problem.

9/18/15

BA

Mentor requested we install a network card for the UPS for remote monitoring.  Network card configured and installed.

6/5/15

BA

On 5 JUN the MGEN sever went down.  On 5 JUN the mentor tried to remotely reboot the MGEN server at I7. It was unsuccessful. On 8 JUN the mentor tried to reboot the server again and it was also unsuccessful. Brandon Androes went to the site to manually reboot and the system came back up.

3/31/15

BA

Mentor said he was receiving data errors from radar.  Cleaned slip rings and contact brushes. Tested cables from slip ring assembly to ACU. All checked out good. RTS at 1815 GMT.  Waiting on feedback from mentor as to whether it is fixed or needs further troubleshooting.

3/13/15

BA

Mentor informed me that CSAPR system showed it was in emergency stop.  Reset emergency stop and interlocks switch at 1700 GMT.

3/5/15

CAM/BA

Mentor advised the servo control was not getting clean data from the elevation motor, similar to what happened a year ago when loose connections were discovered.  Checked elevation cable and found it was still secured with clamps, unbound, and all terminal connections were tight.  Reseated connector at motor and at servo amp.  Error message scrolling on servo amp display: 2 F8022  Enc. 1: enc. signals incorr. (can be cleared in ph. 2).  Did find that the interlock switch for one of the pedestal doors was not engaging properly.   Adjusted the door latch so it now closes tight enough to close the switch.  Notified mentor of findings.

8/6/14

CAM

The input power rocker switch for bus B of the Redundant Ethernet Power Switch (PDU) has failed.  Removed the PDU to the lab for T/S.  Removed the bus B double-pole rocker switch and found the contacts for one pole were stuck open.  Freed the contacts and cleaned.  Reassembled switch and tested.  Switch is now operational and can be used until a replacement is obtained.  Returned PDU to the radar and installed at 2030 GMT.  Powered on and all ports powered up as they should.  Notified mentor that radar was ready for operation.

8/4/14

CAM

Instrument mentor advised the antenna and transmitter control computer (ACU) had lost communications with the transmitter.   T/S found a failed 2-pole (230V) switch on bus B of the DLI Redundant Ethernet Power Switch.  This was preventing power from reaching the power supply that powers the serial to Ethernet adapter between the ACU and transmitter.   No 2-pole (230V) switches available to make repairs.  Mentor will look into sourcing a replacement Ethernet Power Switch.  Site Ops will also order replacement 2-pole switches to repair the old one to keep as a spare.  Radar is offline at this time. 

6/19/14

CAM

The UPS that was originally installed in the system is a 230V input UPS, but site is supplied with 208V, 3-phase power.  230V UPS is faulting frequently due to low input power.  A 208V/230V UPS was ordered to replace the original unit and it has arrived.  R/R *UPS and **battery pack at 1530 GMT.

*Old – JS1028016920   *New – AS1404234228   

**Old – QS1022350479     **New - 7A1410L19263

5/19/14

BC

C-Band UPS fault.  At 1300 GMT noticed a UPS fault on the system. Reset the fault and notified mentor. Mentor said site was operating again at 1850 GMT.

5/14/14

BC

Signal generator was not accessible over network.  Arrived at site at 1300 GMT. Signal generator was configured for the wrong IP address. Programed the IP address and tested connection. Discovered door interlock fault. inspected all interlocks and discovered interlock on back of transmitter rack tripped.  Corrected discrepancy and returned to service 1900 GMT.

5/13/14

BC

UPS faulted.  UPS was faulted showing low input voltage around 1530 according to the radar display. Reset UPS and returned system to service at 2000 GMT.

4/24/14

BC

C-band horizontal data incorrect.  Sent mentor e-mail asking about position of waveguide switches. He said they should be oriented so that the system transmits on both waveguide paths. The system was set so that the waveguide selection was to the tape marked H+V. This position resulted in Vertical only transmission. Changed switch to HV stenciled position and verified waveguide switches were set for dual polarization.

4/24/14

BC

C-band horizontal data incorrect.   Mentor sent an email stating that the LNA in the horizontal channel needed to be replaced. Found spare LNA at the warehouse. At 1430 GMT took unit out of service. Replaced *LNA and RTS 1600 GMT.

*Old - 1565232  *New - 1612466

4/10/14

BC

Signal generator needed calibration.   Removed signal generator from system for calibration at 1400 GMT.

4/8/14

BC

Mentor could not communicate with system.  Tried to view system status remotely and could not reach or ping system. Arrived at site and discovered UPS had a flashing overload light. reset UPS and system rebooted.

1/29/14

CAM

Radar has been encountering intermittent elevation errors.  Checked continuity from elevation motor signal cable to slip ring and found no problems.  Checked connections on opposite side of slip ring and found the nut on terminal 10 was loose, which could possibly cause intermittent problems.  Terminal corresponds to pin 3 (Sin/S2) of elevation motor.  Tightened loose nut and checked all others.  Also found the signal cable 15-pin connector was somewhat loose inside the electronics rack.  Tightened the cable’s screw lock.  Notified mentor of findings and he will run the antenna through some tests when he returns from travel tomorrow.  RTS at 2000 GMT.

10/21/13

CAM

Instrument mentor advised C-band radar had been encountering frequent elevation errors.  Radar was not currently transmitting, but powered azimuth and elevation servos off at 1830 GMT for radome access.  Checked elevation cables and found a cable clamp had come loose, allowing the cables to rub on the elevation gear and also rub on the pedestal as the antenna would spin.  The cables were not physically damaged, but the rubbing may have introduced noise or possibly even snagged just enough at times to cause the errors.  Secured the cable clamp in a location so that the cables could not rub on anything.   Made a rough measurement of the cables at mentor’s request, approximately 14 feet long. Powered servos on and notified mentor of findings.  RTS at 2000 GMT.

10/2/13

CAM

Instrument mentor advised he was unable to access MGEN computer.  Found computer was powered on but unresponsive and nothing displayed on monitor.  Power cycled computer at 1430 GMT.  Computer booted up normally.  RTS at 1445 GMT and notified mentor.

8/6/13

CAM

Instrument mentor advised he was unable to access PGEN computer.  Found computer was powered on but unresponsive and nothing displayed on monitor.  Power cycled computer at 1315 GMT.  Computer booted up normally.  RTS at 1325 GMT and notified mentor.

3/5/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm that resulted in widespread damage to power poles and transmission lines.  Power is now restored but mentor advised MGEN computer hung up on restart.  Traveled to site and power cycled MGEN computer.  Contacted mentor and verified communications were restored.  RTS at 1800 GMT.

12/24/12

CAM

Instrument mentor requested the transmitter interface box be removed and sent to Chad Baldi at ProSensing.  The interface will be taken to Manus to be installed and the failed unit from Manus will be repaired and returned to SGP for use here.  Removed transmitter interface box at 1545 GMT and delivered to S/R for shipment.  Radar was already offline due to previous cable failure in radome.   *Old – 1107-01  (Part# TR-1113-PLC, date code 1107)

5/22/12

CAM

Stolhand performed seasonal check of HVAC and window unit.  No problems.

4/10/12

SPD

Wave guide will not hold pressure.  Worked with mentor and discovered that there was a wave guide leak that occurs when the antenna is rotating. Wave guide piece above AZ rotary joint on the "H" channel has a leak in the weld of one of the 90 degree bends. Made a video and sent it to the mentor. Awaiting further instructions from the mentor. RTS @ 1815 GMT.

4/9/12

SPD

Found the Dehydrator inoperative with alarm code A01 and A11. Cycled the Dehydrator power at 1410 GMT and it started to pressure the waveguide. Wave guide pressured to 5.8 psi.  The transmitter came on but it sounded like there was arcing inside the waveguide. Placed the transmitter into local and standby at 1434 GMT to check for waveguide leaks and tighten the waveguide screws. Left the transmitter in remote and standby.  Notified mentor and suggest having someone at the site before mentor tries to remotely restart the RF.

4/7/12

Mentor

Mentor placed C-SAPR back into operation.  Will need to check for leaks on Monday.

4/5/12

SPD

At 1450 GMT Dehydrator was not running.

Pressure = .4 psi.

Status = L L .4 (Leaky system (with decimal point if regenerating).

Error codes

A01= Low Pressure alarm (unit cannot attain desired pressure).

A11= Excessive compressor operation (reset by power off).

Power cycled dehydrator unit @ 1455 GMT. Wave guide pressure climbed to 5.9 psi. There were no error codes. We then checked the azimuth rotary joint and surrounding wave guide connections for leaks. None were found.  At 1543 GMT the dehydrator indicated a duty cycle of 6. Returned to the site at 1855 GMT and the dehydrator was still holding with a duty cycle of 6.

So when the antenna is stationary there doesn’t appear to be a leak. When the antenna is returned to service and a leak reoccurs my thought would be a rotary joint problem.   Further troubleshooting using gas barriers or other methods would determine which rotary joint (azimuth or elevation) have the problem.

3/28/12

SPD

C-Band dehydrator had an A09 error code.  Replaced number 2 dryer assembly. C-Band dehydrator RTS @ 1650 GMT.

2/3/12

SD

Received a phone call from Kim Weaver stating that the dehydrator had an alarm and that there must be a air leak and that he had shut down the system.  After proper channels were cleared David Breedlove and I went to the site. Upon inspection it was discovered that the wave guide attached to the upper portion of the AZ rotary joint was loose. Upon tightening the wave guide the leak disappeared. Checking the Dehydrator we had a duty cycle of 11 and dropped to 10 as the system cycled. Kim Weaver stated that was a good number and started the radar system remotely. When Dave and I left the pedestal was spinning and Kim was powering up the transmitter.

2/1/12

SD

Met with Kim Weaver and Jack. Kim was performing a software upgrade while Jack was working on the Radar pedestal. Kim provided helpful information about the Radar system. He could not go into as much detail as he wanted due to the radar being powered off for the upgrade.

 

Jack discussed the pedestal and information about periodic maintenance. Greased the AZ and EL bull gears and lubricated the AZ and EL main bearings. There has been a problem with AZ duel rotary joint wobbling. Adjustment of the EL rotary joints, both the Horizontal and Vertical waveguide runs and some work on the rotary joint mounts reduced the wobble to barely perceptible. After the adjustments the bracket used to hold the AZ Duel Rotary joint in place will no longer line up so it is not in use. We then checked that the waveguide was holding pressure. The dehydrator showed a duty cycle of 2.

 

Kim was still not finished with the software upgrade and it was getting late so I left for the day.

 

Was informed on February 02 2012 that The radar was up and running and all was well.

1/2012

ATSC

The  failed dual rotary joint  at the C-SAPR  was replaced  during the calibration trip.   Waveguide pressurization was then tested and held at 400 mb with a 2% duty cycle.  The antenna was rotated several

times to check for alignment and showed considerable improvement from the previous install.  However, even with the securing bolts removed  that hold the rotary joint in place, a misalignment still exists, which was seen by the slight wobble as the antenna was rotated in azimuth.  The system was left in stand-by until ARC personnel align the waveguide so the current joint does not fail.

9/19/11

DKB

1800 – 2000 GMT.  Mentor notified us via e-mail that the ACU computer would not respond. Found the ACU computer powered off and restarted at 1820 GMT. Vendor was able to log on. Found a waveguide fault and dehumidifier alarm. Cycled the power to the Dehumidifier to reset. Vendor indicated that the antenna could not find its azimuth reference zero position. Cycled the power to the azimuth drive several times did not correct. Left azimuth inoperative awaiting vendor visit.